# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ 9 Limitations and future work
Current implementations may require more training iterations than MLPs for certain tasks, though the per-iteration computational cost is typically lower due to fewer parameters. When memory-efficient sequential computation is employed to enable training of wider architectures, wall-clock training time increases, representing a fundamental trade-off between memory usage and computational efficiency. The sequential computation mode proves most beneficial for architectures with individual layers exceeding 128-256 units in width, while offering minimal advantage for very deep networks with modest layer widths where the memory bottleneck lies in storing activations across many layers rather than within-block computations. The development of adaptive knot placement strategies that concentrate resolution where data lives while maintaining fixed parameter counts could improve both efficiency and interpretability.
The lateral mixing mechanism, while empirically beneficial, lacks theoretical justification within the Sprecher framework. Understanding whether this enhancement can be connected to the underlying mathematical structure or represents a purely empirical improvement remains an open question. Future work could explore adaptive mixing topologies beyond cyclic patterns, potentially learning the neighborhood structure $\mathcal{N}(q)$ itself, or investigating connections to graph neural networks where the mixing pattern could be viewed as a learnable graph structure over output dimensions.
@@
Architectural enhancements also merit exploration. Unlike KANs where edge pruning can significantly reduce parameters, SNs’ vector structure suggests different optimization strategies. Automatic pruning of entire Sprecher blocks based on their contribution to network output could yield more compact architectures. The learned lateral mixing weights could guide this pruning: blocks with near-zero mixing weights might be candidates for removal. For very high-dimensional problems where even $O(N)$ scaling becomes prohibitive, hybrid approaches using low-rank approximations for $\lambda^{(\ell)}$ or replacing splines with small neural sub-networks in certain blocks could maintain efficiency while improving expressivity. Alternatively, hierarchical lateral mixing patterns (e.g., mixing within local groups before global mixing) could provide a middle ground between full connectivity and the current nearest-neighbor approach.
@@ 10 Conclusion
The key contributions are demonstrating that Sprecher’s shallow, highly-structured formula can be extended into an effective deep architecture with remarkable parameter efficiency: $O(LN+LG)$ compared to MLPs’ $O(LN^{2})$ or KANs’ $O(LN^{2}G)$, and enabling $O(LN)$ forward working memory (peak intermediate storage) through sequential computation strategies compared to MLPs’ $O(LN^{2})$. This dual efficiency in both parameters and memory comes from adhering to Sprecher’s use of weight vectors rather than matrices, representing a strong architectural constraint that may serve as either beneficial inductive bias or limitation depending on the problem domain. The addition of lateral mixing connections provides a parameter-efficient mechanism for intra-block communication, partially addressing the limitations of the constrained weight structure while maintaining the overall efficiency of the architecture.
Our initial demonstrations show SNs can successfully learn diverse functions and achieve competitive performance across synthetic regression benchmarks and PDE-constrained learning problems, with the added benefit of interpretable spline visualizations. The lateral mixing mechanism proves particularly valuable for vector-valued outputs and deeper networks, often improving optimization and convergence behavior even though the learned $\Phi$ splines can remain highly oscillatory. The sequential computation mode enables training of architectures with wide layers that would otherwise exhaust available memory, making SNs particularly suitable for exploring high-capacity models under memory constraints. However, the need for more training iterations in some cases and theoretical gaps regarding deep network universality remain open questions. The theoretical status of lateral mixing — whether it represents a principled extension of Sprecher’s construction or merely an empirical enhancement — requires further investigation.
Whether SNs prove broadly useful or remain a fascinating special case, they demonstrate the value of mining classical mathematical results for architectural inspiration in modern deep learning. The successful integration of lateral mixing shows how theorem-inspired designs can be enhanced with empirical innovations while maintaining their core theoretical structure. The ability to evaluate SN layers with $O(LN)$ forward working memory (peak intermediate storage) distinguishes SNs from other major architectures, suggesting applications in memory-constrained settings and when exploring extremely wide output layers. This synthesis of rigorous mathematical foundations with practical deep learning techniques points toward a promising direction for developing novel architectures that balance theoretical elegance with empirical effectiveness.

### v2

@@ 9 Limitations and future work
Current implementations may require more training iterations than MLPs for certain tasks; per-iteration wall-clock cost is workload- and implementation-dependent (fewer parameters reduce optimizer overhead, but spline evaluation and memory-saving recomputation can increase runtime). When memory-efficient sequential computation is employed to enable training of wider architectures, wall-clock training time increases, representing a fundamental trade-off between memory usage and computational efficiency. The sequential computation mode proves most beneficial for architectures with individual layers exceeding 128-256 units in width, while offering minimal advantage for very deep networks with modest layer widths where the memory bottleneck lies in storing activations across many layers rather than within-block computations. The development of adaptive knot placement strategies that concentrate resolution where data lives while maintaining fixed parameter counts could improve both efficiency and interpretability.
The lateral mixing mechanism, while empirically beneficial, lacks theoretical justification within the Sprecher framework. Understanding whether this enhancement can be connected to the underlying mathematical structure or represents a purely empirical improvement remains an open question. Future work could explore adaptive mixing topologies beyond cyclic patterns, potentially learning the neighborhood structure $\mathcal{N}(q)$ itself, or investigating connections to graph neural networks where the mixing pattern could be viewed as a learnable graph structure over output dimensions. Another promising direction is additional sharing across depth, e.g., tying $\Phi^{(\ell)}$ across groups of layers or learning a small dictionary of outer splines that multiple blocks reuse, which could further reduce parameters and potentially improve interpretability.
@@
Architectural enhancements also merit exploration. Unlike KANs where edge pruning can significantly reduce parameters, SNs’ vector structure suggests different optimization strategies. Automatic pruning of entire Sprecher blocks based on their contribution to network output could yield more compact architectures. The learned lateral mixing weights could guide this pruning: blocks with near-zero mixing weights might be candidates for removal. Preliminary observations suggest that in deep networks, $\Phi$ splines in adjacent blocks sometimes converge to similar shapes, raising the question of whether certain splines could be shared across blocks or removed entirely for further parameter reduction. Investigating when and why this similarity emerges—and whether it can be exploited via explicit cross-block spline tying or regularization encouraging such sharing—could yield more compact architectures without sacrificing accuracy. For very high-dimensional problems where even $O(N)$ scaling becomes prohibitive, hybrid approaches using low-rank approximations for $\lambda^{(\ell)}$ or replacing splines with small neural sub-networks in certain blocks could maintain efficiency while improving expressivity. Alternatively, hierarchical lateral mixing patterns (e.g., mixing within local groups before global mixing) could provide a middle ground between full connectivity and the current nearest-neighbor approach.
@@ 10 Conclusion
The key contributions are demonstrating that Sprecher’s shallow, highly-structured formula can be extended into an effective deep architecture with remarkable parameter efficiency: $O(LN+LG)$ compared to MLPs’ $O(LN^{2})$ or KANs’ $O(LN^{2}G)$, and enabling $O(LN)$ forward working memory (parameters + peak forward intermediates) through sequential computation strategies compared to MLPs’ $O(LN^{2})$. This dual efficiency in both parameters and memory comes from adhering to Sprecher’s use of weight vectors rather than matrices, representing a strong architectural constraint that may serve as either a beneficial inductive bias or a limitation depending on the problem domain. The linear memory scaling is particularly relevant for deployment on resource-constrained devices such as embedded sensors, microcontrollers, and edge computing platforms where memory is severely limited and expensive. It may also benefit environmentally-conscious deployment scenarios where reducing memory requirements translates to lower power consumption. The addition of lateral mixing connections provides a parameter-efficient mechanism for intra-block communication, partially addressing the limitations of the constrained weight structure while maintaining the overall efficiency of the architecture.
Our initial demonstrations show SNs can successfully learn diverse functions and achieve competitive performance across synthetic regression benchmarks and PDE-constrained learning problems, with the added benefit of interpretable spline visualizations. The lateral mixing mechanism proves particularly valuable for vector-valued outputs and deeper networks, often improving optimization and convergence behavior even though the learned $\Phi$ splines may exhibit complex, non-smooth structure. The sequential computation mode enables training of architectures with wide layers that would otherwise exhaust available memory, making SNs particularly suitable for exploring high-capacity models under memory constraints. However, the need for more training iterations in some cases and theoretical gaps regarding deep network universality remain open questions. The theoretical status of lateral mixing — whether it represents a principled extension of Sprecher’s construction or merely an empirical enhancement — requires further investigation.
Whether SNs prove broadly useful or remain a fascinating special case, they demonstrate the value of mining classical mathematical results for architectural inspiration in modern deep learning. The successful integration of lateral mixing shows how theorem-inspired designs can be enhanced with empirical innovations while maintaining their core theoretical structure. The ability to evaluate SN layers with $O(LN)$ forward working memory distinguishes SNs from other major architectures, suggesting applications in memory-constrained settings and when exploring extremely wide output layers. This synthesis of rigorous mathematical foundations with practical deep learning techniques points toward a promising direction for developing novel architectures that balance theoretical elegance with empirical effectiveness.